# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ D.3 Corollary 1
$\max_{n+1\leq i\leq n+m}\|{\hat{Z}_{i}-Z_{i}}\|=O_{P}(\frac{1}{m^{1/2-c^{\prime}}})$
@@
$\max_{n+1\leq i\leq n+m}\|{\hat{Z}_{i}-Z_{i}}\|=O_{P}(\frac{(\log m)^{c^{\prime}}}{\sqrt{m}})$
@@
By applying Theorem 4, with $\delta_{m}=\frac{1}{m^{1/2-c^{\prime}}}$ and $\frac{(\log m)^{c^{\prime}}}{\sqrt{m}}$, respectively. We can see that Assumption A4 holds.
@@ Appendix E Proof of Theorem 5 (Convergence of Network Moments)
Let $X_{1},X_{2},\dotsc,X_{n}$ be independent and identically distributed (i.i.d.) random variables from a continuous distribution $F$ on ${\bf R}^{d}$. Suppose we have approximate observations $\hat{X}_{1},\hat{X}_{2},\dotsc,\hat{X}_{n}$, which may be dependent, but whose empirical distribution function $\hat{F}_{n}$ converges uniformly in probability to $F$. We are interested in understanding the relationship between U-statistics computed from the exact data and those computed from the approximate data.
Let $h:({\bf R}^{d})^{r}\to{\bf R}$ be a bounded, Lipchitz continuous, and symmetric kernel function. We define the U-statistics:
@@
where $\delta_{(X_{i_{1}},\dots,X_{i_{r}})}$ is the Dirac measure at the point $(X_{i_{1}},\dots,X_{i_{r}})$ in $({\bf R}^{d})^{r}$.
@@
- The product measure of $F$ is $F^{\otimes r}$, defined on $({\bf R}^{d})^{r}$ as:
$F^{\otimes r}(A)=\int_{A}dF(x_{1})\cdots dF(x_{r}),$
for measurable sets $A\subset({\bf R}^{d})^{r}$.
@@
Let $F$ be a CDF on ${\bf R}^{d}$ and $\hat{F}_{n}$ be empirical distribution functions on ${\bf R}^{d}$ defined as:
@@
    $\sup_{x\in{\bf R}^{d}}|\hat{F}_{n}(x)-F(x)|\xrightarrow{P}0\quad\text{as }n\to\infty.$
Then, for each $m\geq 1$, the product measures $F_{n}^{\otimes r}$ and $\hat{F}_{n}^{\otimes r}$ converge weakly to $F^{\otimes r}$ on $({\bf R}^{d})^{r}$, that is:
$\hat{F}_{n}^{\otimes r}$ $\xrightarrow{\text{w.P.}}F^{\otimes r},$
where $\xrightarrow{\text{w.P.}}$ denotes weak convergence in probability.
@@
Let $F_{n}$ and $\hat{F}_{n}$ be empirical distribution functions on ${\bf R}^{d}$ defined as:
@@
  • $X_{1},X_{2},\dotsc,X_{n}$ are independent and identically distributed (i.i.d.) random variables with distribution $F$ on ${\bf R}^{d}$.
@@
    $\sup_{x\in{\bf R}^{d}}|\hat{F}_{n}(x)-F(x)|\xrightarrow{P}0\quad\text{as }n\to\infty.$
Then, for each $m\geq 1$, the product measures $F_{n}^{\otimes r}$ and $\hat{F}_{n}^{\otimes r}$ converge weakly to $F^{\otimes r}$ on $({\bf R}^{d})^{r}$, that is:
$F_{n}^{\otimes r}$ $\xrightarrow{\text{w.a.s.}}F^{\otimes r},$
$\hat{F}_{n}^{\otimes r}$ $\xrightarrow{\text{w.P.}}F^{\otimes r},$
where $\xrightarrow{\text{w.a.s.}}$ denotes weak convergence almost surely, and $\xrightarrow{\text{w.P.}}$ denotes weak convergence in probability.
@@
$\sup_{x\in{\bf R}^{d}}|F_{n}(x)-F(x)|\xrightarrow{\text{a.s.}}0\quad\text{as }n\to\infty.$
This implies that for any bounded, continuous function $g:{\bf R}^{d}\to{\bf R}$,
$\int g\,dF_{n}\xrightarrow{\text{a.s.}}\int g\,dF.$
Now, consider any bounded, continuous function $h:({\bf R}^{d})^{r}\to{\bf R}$. We need to show that:
$\int h\,dF_{n}^{\otimes r}\xrightarrow{\text{a.s.}}\int h\,dF^{\otimes r}.$
@@
$\int h(x_{1},x_{2},\dotsc,x_{r})\,dF_{n}^{\otimes r}(x_{1},x_{2},\dotsc,x_{r})=\int\cdots\int h(x_{1},x_{2},\dotsc,x_{r})\,dF_{n}(x_{1})\cdots dF_{n}(x_{r}).$
@@
$\int h(x_{1},x_{2},\dotsc,x_{r})\,dF^{\otimes r}(x_{1},x_{2},\dotsc,x_{r})=\int\cdots\int h(x_{1},x_{2},\dotsc,x_{r})\,dF(x_{1})\cdots dF(x_{r}).$
Since $F_{n}\xrightarrow{\text{a.s.}}F$ uniformly, for each $j=1,2,\dotsc,m$, we have:
$\int g_{j}\,dF_{n}\xrightarrow{\text{a.s.}}\int g_{j}\,dF,$
for any bounded, continuous function $g_{j}:{\bf R}^{d}\to{\bf R}$.
@@
$\int h\,dF_{n}^{\otimes r}\xrightarrow{\text{a.s.}}\int h\,dF^{\otimes r}.$
@@
$\hat{U}_{n}-U_{n}\xrightarrow{P}0\quad\text{as }n\to\infty.$
@@
We aim to show that $\hat{U}_{n}-U_{n}\xrightarrow{P}0$ as $n\to\infty$. We will use
$\theta=\int h\,dF^{\otimes r}=\mathbb{E}[h(X_{1},\dotsc,x_{r})]$
as the intermediate quantity for the convergence. To directly characterize the approximation error between the U-statistic and the integral over the product measure, we consider the difference between sampling without replacement (as in $U_{n},\hat{U}_{n}$) and sampling with replacement (as in $\int h\,d{F}_{n}^{\otimes r},\int h\,d\hat{F}_{n}^{\otimes r}$).
@@
$S_{\text{distinct}}=\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}h(\hat{X}_{i_{1}},\dotsc,\hat{X}_{i_{r}})$
@@
$S_{\text{all}}=\sum_{i_{1}=1}^{n}\cdots\sum_{i_{r}=1}^{n}h(\hat{X}_{i_{1}},\dotsc,\hat{X}_{i_{r}}),$
@@
$\hat{U}_{n}=\frac{S_{\text{distinct}}}{\binom{n}{r}},\quad\int h\,d\hat{F}_{n}^{\otimes r}=\frac{S_{\text{all}}}{n^{r}}.$
@@
$\varepsilon_{n}=\hat{U}_{n}-\int h\,d\hat{F}_{n}^{\otimes r}=\frac{S_{\text{distinct}}}{\binom{n}{r}}-\frac{S_{\text{all}}}{n^{r}}.$
@@
$S_{\text{all}}=r!S_{\text{distinct}}+R_{n},$
@@
$\binom{n}{r}r!=n^{r}\left(1-O\left(\frac{1}{n}\right)\right),\quad n^{r}-\binom{n}{r}r!=n^{r}O\left(\frac{1}{n}\right)=O(n^{r-1}).$
From this, we get $\hat{U}_{n}=\frac{S_{\text{distinct}}}{\binom{n}{r}}=\frac{S_{\text{all}}-R_{n}}{r!\binom{n}{r}}$. Compare this with $\int h\,d\hat{F}_{n}^{\otimes r}$, we have
$\hat{U}_{n}-\int h\,d\hat{F}_{n}^{\otimes r}=\frac{S_{\text{all}}-R_{n}}{r!\binom{n}{r}}-\frac{S_{\text{all}}}{n^{r}}=S_{\text{all}}\left(\frac{1}{r!\binom{n}{r}}-\frac{1}{n^{r}}\right)-\frac{R_{n}}{r!\binom{n}{r}}$.
With the boundedness assumption of $h$, we have $S_{\text{all}}\leq Mn^{r}$, thus:
$S_{\text{all}}\left(\frac{1}{r!\binom{n}{r}}-\frac{1}{n^{r}}\right)=Mn^{r}O\left(\frac{1}{n^{r+1}}\right)=O\left(\frac{1}{n}\right).$
Similarly, we have $|R_{n}|=MO(n^{r-1})$ and $\frac{1}{r!\binom{n}{r}}=O\left(\frac{1}{n}\right).$ Combining these leads to
$|\hat{U}_{n}-\int h\,d\hat{F}_{n}^{\otimes r}|\leq O\left(\frac{1}{n}\right)+O\left(\frac{1}{n}\right)=O\left(\frac{1}{n}\right).$
Thus $\hat{U}_{n}-\int h\,d\hat{F}_{n}^{\otimes r}=O(1/n)\to 0$ as $n\to\infty$.
Using the same reasoning, define $\varepsilon_{n}={U}_{n}-\int h\,d{F}_{n}^{\otimes r}$ and we also have $|{\varepsilon}_{n}|=O(n^{-1}).$
@@
$\hat{U}_{n}-U_{n}$ $=\left(\hat{U}_{n}-\int h\,d\hat{F}_{n}^{\otimes r}\right)+\left(\int h\,d\hat{F}_{n}^{\otimes r}-\theta\right)+\left(\theta-\int h\,dF_{n}^{\otimes r}\right)+\left(\int h\,dF_{n}^{\otimes r}-U_{n}\right)$
$=\hat{\varepsilon}_{n}+\left(\int h\,d\hat{F}_{n}^{\otimes r}-\theta\right)+\left(\theta-\int h\,dF_{n}^{\otimes r}\right)-\varepsilon_{n}.$
@@
$|\hat{U}_{n}-U_{n}|\leq|\hat{\varepsilon}_{n}|+\left|\int h\,d\hat{F}_{n}^{\otimes r}-\int h\,dF^{\otimes r}\right|+\left|\int h\,dF^{\otimes r}-\int h\,dF_{n}^{\otimes r}\right|+|\varepsilon_{n}|.$
@@
$\hat{U}_{n}-U_{n}\xrightarrow{P}0\quad\text{as }n\to\infty.$
@@
where $A_{[i_{1},i_{2},\ldots,i_{r}]}$ denotes the induced subgraph on vertices $\{i_{1},\ldots,i_{r}\}$. For every $\varepsilon>0$, there exist positive constants $c(r)$ (not depending on $n$) such that
$\mathbb{P}\bigl{(}|X-\mathbb{E}[X]|\geq\varepsilon\bigr{)}\leq 2\exp(-c(r)\varepsilon^{2}n^{2}),$
@@
$X\xrightarrow{p}\mathbb{E}[X]\quad\text{as }n\to\infty.$
@@
Since $\sum_{j=1}^{r}\Delta_{j}^{2}=O(1/n^{2})$, we have
@@
$X\xrightarrow{p}\mathbb{E}[X]\quad\text{as }n\to\infty.$
@@
$U_{n}=\mathbb{E}\big{[}X_{R}(A)|\{Z_{i}\}\big{]}=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}\mathbb{E}\big{[}{\mathbb{I}}(A_{[i_{1},\cdots,i_{r}]}\cong R)|\{Z_{i}\}\big{]}=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}h(Z_{i_{1}},\cdots,Z_{i_{r}})$
@@
$\hat{U}_{n}=\mathbb{E}\big{[}X_{R}(A)|\{\hat{Z}_{i}\}\big{]}=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}h(\hat{Z}_{i_{1}},\cdots,\hat{Z}_{i_{r}}).$
@@
$).\displaystyle X_{R}(A)-X_{R}(\tilde{A})=\big{(}X_{R}(A)-U_{n}\big{)}+\big{(}U_{n}-\hat{U}_{n}\big{)}+\big{(}\hat{U}_{n}-X_{R}(\tilde{A}$$\big{)}$.
Conditioning on $\{Z_{i}\}$ and $\{\hat{Z}_{i}\}$, respectively, the first and third term vanish in probability from Lemma E.4. By Theorem 3, we know that the empirical CDF of $\tilde{Z}_{i}$’s satisfies the requirement of Lemma E.2. And then we call Lemma E.3, and the second term also vanishes in probability. ∎

### v3

@@ D.3 Corollary 1
$\max_{n+1\leq i\leq n+m}\|{\hat{Z}_{i}-Z_{i}}\|=\mathcal{O}_{\mathbb{P}}\left(\frac{1}{m^{1/2-c^{\prime}}}\right)$
@@
$\max_{n+1\leq i\leq n+m}\|{\hat{Z}_{i}-Z_{i}}\|=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log m)^{c^{\prime}}}{\sqrt{m}}\right)$
@@
By applying Theorem 4, with $\delta_{m}=1/m^{1/2-c^{\prime}}$ and $(\log m)^{c^{\prime}}/\sqrt{m}$, respectively. We can see that Assumption A4 holds.
@@ Appendix E Proof of Theorem 5 (Convergence of Network Moments)
Let $X_{1},X_{2},\dotsc,X_{n}$ be independent and identically distributed (i.i.d.) random variables from a continuous distribution $F$ on $\mathbb{R}^{d}$. Suppose we have approximate observations $\hat{X}_{1},\hat{X}_{2},\ldots,\hat{X}_{n}$, which may be dependent, but whose empirical distribution function $\hat{F}_{n}$ converges uniformly in probability to $F$. We are interested in understanding the relationship between U-statistics computed from the exact data and those computed from the approximate data.
Let $h:(\mathbb{R}^{d})^{r}\to\mathbb{R}$ be a bounded, Lipchitz continuous, and symmetric kernel function. We define the U-statistics:
@@
where $\delta_{(X_{i_{1}},\dots,X_{i_{r}})}$ is the Dirac measure at the point $(X_{i_{1}},\dots,X_{i_{r}})$ in $(\mathbb{R}^{d})^{r}$.
@@
- The product measure of $F$ is $F^{\otimes r}$, defined on $(\mathbb{R}^{d})^{r}$ as:
$F^{\otimes r}(A)=\int_{\mathcal{A}}\,\mathrm{d}F(x_{1})\cdots\mathrm{d}F(x_{r}),$
for measurable sets $\mathcal{A}\subset(\mathbb{R}^{d})^{r}$.
@@
Let $F$ be a CDF on $\mathbb{R}^{d}$ and $\hat{F}_{n}$ be empirical distribution functions on $\mathbb{R}^{d}$ defined as:
@@
  • $\hat{X}_{1},\hat{X}_{2},\ldots,\hat{X}_{n}$ are random variables (not necessarily independent) such that:
    $\sup_{x\in\mathbb{R}^{d}}\big|\hat{F}_{n}(x)-F(x)\big|\xrightarrow{\mathbb{P}}0\quad\text{as }n\to\infty.$
Then, for each $m\geq 1$, the product measures $F_{n}^{\otimes r}$ and $\hat{F}_{n}^{\otimes r}$ converge weakly to $F^{\otimes r}$ on $(\mathbb{R}^{d})^{r}$, that is:
$\hat{F}_{n}^{\otimes r}$ $\xrightarrow{\mathrm{w.}\mathbb{P}}F^{\otimes r},$
where $\xrightarrow{\mathrm{w.}\mathbb{P}}$ denotes weak convergence in probability.
@@
Let $F_{n}$ and $\hat{F}_{n}$ be empirical distribution functions on $\mathbb{R}^{d}$ defined as:
@@
  • $X_{1},X_{2},\dotsc,X_{n}$ are independent and identically distributed (i.i.d.) random variables with distribution $F$ on $\mathbb{R}^{d}$.
@@
    $\sup_{x\in\mathbb{R}^{d}}\big|\hat{F}_{n}(x)-F(x)\big|\xrightarrow{\mathbb{P}}0\quad\text{as }n\to\infty.$
Then, for each $m\geq 1$, the product measures $F_{n}^{\otimes r}$ and $\hat{F}_{n}^{\otimes r}$ converge weakly to $F^{\otimes r}$ on $(\mathbb{R}^{d})^{r}$, that is:
$F_{n}^{\otimes r}\xrightarrow{\mathrm{w.a.s.}}F^{\otimes r},\quad\hat{F}_{n}^{\otimes r}\xrightarrow{\mathrm{w.\mathbb{P}}}F^{\otimes r},$
where $\xrightarrow{\mathrm{w.a.s.}}$ denotes weak convergence almost surely, and $\xrightarrow{\mathrm{w.\mathbb{P}}}$ denotes weak convergence in probability.
@@
$\sup_{x\in\mathbb{R}^{d}}\big|F_{n}(x)-F(x)\big|\xrightarrow{\mathrm{a.s.}}0\quad\text{as }n\to\infty.$
This implies that for any bounded, continuous function $g:\mathbb{R}^{d}\to\mathbb{R}$,
$\int g\,\mathrm{d}F_{n}\xrightarrow{\mathrm{a.s.}}\int g\,\mathrm{d}F.$
Now, consider any bounded, continuous function $h:(\mathbb{R}^{d})^{r}\to\mathbb{R}$. We need to show that:
$\int h\,\mathrm{d}F_{n}^{\otimes r}\xrightarrow{\mathrm{a.s.}}\int h\,\mathrm{d}F^{\otimes r}.$
@@
$\int h(x_{1},x_{2},\dotsc,x_{r})\,\mathrm{d}F_{n}^{\otimes r}(x_{1},x_{2},\dotsc,x_{r})=\int\cdots\int h(x_{1},x_{2},\dotsc,x_{r})\,\mathrm{d}F_{n}(x_{1})\cdots\mathrm{d}F_{n}(x_{r}).$
@@
$\int h(x_{1},x_{2},\dotsc,x_{r})\,\mathrm{d}F^{\otimes r}(x_{1},x_{2},\dotsc,x_{r})=\int\cdots\int h(x_{1},x_{2},\dotsc,x_{r})\,\mathrm{d}F(x_{1})\cdots\mathrm{d}F(x_{r}).$
Since $F_{n}\xrightarrow{\mathrm{a.s.}}F$ uniformly, for each $j=1,2,\dotsc,m$, we have:
$\int g_{j}\,\mathrm{d}F_{n}\xrightarrow{\mathrm{a.s.}}\int g_{j}\,\mathrm{d}F$
for any bounded, continuous function $g_{j}:\mathbb{R}^{d}\to\mathbb{R}$.
@@
$\int h\,\mathrm{d}F_{n}^{\otimes r}\xrightarrow{\mathrm{a.s.}}\int h\,\mathrm{d}F^{\otimes r}.$
@@
$\hat{U}_{n}-U_{n}\xrightarrow{\mathbb{P}}0\quad\text{as }n\to\infty.$
@@
We aim to show that $\hat{U}_{n}-U_{n}\xrightarrow{\mathbb{P}}0$ as $n\to\infty$. We will use
$\theta=\int h\,\mathrm{d}F^{\otimes r}=\mathbb{E}[h(X_{1},\dotsc,x_{r})]$
as the intermediate quantity for the convergence. To directly characterize the approximation error between the U-statistic and the integral over the product measure, we consider the difference between sampling without replacement (as in $U_{n},\hat{U}_{n}$) and sampling with replacement (as in $\int h\,\mathrm{d}{F}_{n}^{\otimes r},\int h\,\mathrm{d}\hat{F}_{n}^{\otimes r}$).
@@
$S_{\mathrm{distinct}}=\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}h(\hat{X}_{i_{1}},\dotsc,\hat{X}_{i_{r}})$
@@
$S_{\mathrm{all}}=\sum_{i_{1}=1}^{n}\cdots\sum_{i_{r}=1}^{n}h(\hat{X}_{i_{1}},\dotsc,\hat{X}_{i_{r}}),$
@@
$\hat{U}_{n}=\frac{S_{\mathrm{distinct}}}{\binom{n}{r}},\quad\int h\,\mathrm{d}\hat{F}_{n}^{\otimes r}=\frac{S_{\mathrm{all}}}{n^{r}}.$
@@
$\varepsilon_{n}=\hat{U}_{n}-\int h\,\mathrm{d}\hat{F}_{n}^{\otimes r}=\frac{S_{\mathrm{distinct}}}{\binom{n}{r}}-\frac{S_{\mathrm{all}}}{n^{r}}.$
@@
$S_{\mathrm{all}}=r!S_{\mathrm{distinct}}+R_{n},$
@@
$\binom{n}{r}r!=n^{r}\left(1-\mathcal{O}\left(\frac{1}{n}\right)\right),\quad n^{r}-\binom{n}{r}r!=n^{r}\mathcal{O}\left(\frac{1}{n}\right)=\mathcal{O}(n^{r-1}).$
From this, we get $\hat{U}_{n}=S_{\mathrm{distinct}}/\binom{n}{r}=(S_{\mathrm{all}}-R_{n})/(r!\binom{n}{r})$. Compare this with $\int h\,\mathrm{d}\hat{F}_{n}^{\otimes r}$, we have
$\hat{U}_{n}-\int h\,\mathrm{d}\hat{F}_{n}^{\otimes r}=\frac{S_{\mathrm{all}}-R_{n}}{r!\binom{n}{r}}-\frac{S_{\mathrm{all}}}{n^{r}}=S_{\mathrm{all}}\left(\frac{1}{r!\binom{n}{r}}-\frac{1}{n^{r}}\right)-\frac{R_{n}}{r!\binom{n}{r}}$.
With the boundedness assumption of $h$, we have $S_{\mathrm{all}}\leq Mn^{r}$, thus:
$S_{\mathrm{all}}\left(\frac{1}{r!\binom{n}{r}}-\frac{1}{n^{r}}\right)=Mn^{r}\mathcal{O}\left(\frac{1}{n^{r+1}}\right)=\mathcal{O}\left(\frac{1}{n}\right).$
Similarly, we have $|R_{n}|=M\mathcal{O}(n^{r-1})$ and $1/(r!\binom{n}{r})=\mathcal{O}(1/n).$ Combining these leads to
$\left|\hat{U}_{n}-\int h\,\mathrm{d}\hat{F}_{n}^{\otimes r}\right|\leq\mathcal{O}\left(\frac{1}{n}\right)+\mathcal{O}\left(\frac{1}{n}\right)=\mathcal{O}\left(\frac{1}{n}\right).$
Thus $\hat{U}_{n}-\int h\,\mathrm{d}\hat{F}_{n}^{\otimes r}=\mathcal{O}(1/n)\to 0$ as $n\to\infty$.
Using the same reasoning, define $\varepsilon_{n}={U}_{n}-\int h\,\mathrm{d}{F}_{n}^{\otimes r}$ and we also have $|{\varepsilon}_{n}|=\mathcal{O}(1/n).$
@@
$\hat{U}_{n}-U_{n}$ $=\left(\hat{U}_{n}-\int h\,\mathrm{d}\hat{F}_{n}^{\otimes r}\right)+\left(\int h\,\mathrm{d}\hat{F}_{n}^{\otimes r}-\theta\right)+\left(\theta-\int h\,\mathrm{d}F_{n}^{\otimes r}\right)+\left(\int h\,\mathrm{d}F_{n}^{\otimes r}-U_{n}\right)$
$=\hat{\varepsilon}_{n}+\left(\int h\,\mathrm{d}\hat{F}_{n}^{\otimes r}-\theta\right)+\left(\theta-\int h\,\mathrm{d}F_{n}^{\otimes r}\right)-\varepsilon_{n}.$
@@
$|\hat{U}_{n}-U_{n}|\leq|\hat{\varepsilon}_{n}|+\left|\int h\,\mathrm{d}\hat{F}_{n}^{\otimes r}-\int h\,\mathrm{d}F^{\otimes r}\right|+\left|\int h\,\mathrm{d}F^{\otimes r}-\int h\,\mathrm{d}F_{n}^{\otimes r}\right|+|\varepsilon_{n}|.$
@@
$\hat{U}_{n}-U_{n}\xrightarrow{\mathbb{P}}0\quad\text{as }n\to\infty.$
@@
where $A_{[i_{1},i_{2},\ldots,i_{r}]}$ denotes the induced subgraph on vertices $\{i_{1},i_{2},\ldots,i_{r}\}$. For every $\varepsilon>0$, there exist positive constants $c(r)$ (not depending on $n$) such that
$\mathbb{P}(|X-\mathbb{E}[X]|\geq\varepsilon)\leq 2\exp(-c(r)\varepsilon^{2}n^{2}),$
@@
$X\xrightarrow{\mathbb{P}}\mathbb{E}[X]\quad\text{as }n\to\infty.$
@@
Since $\sum_{j=1}^{r}\Delta_{j}^{2}=\mathcal{O}(1/n^{2})$, we have
@@
$X\xrightarrow{\mathbb{P}}\mathbb{E}[X]\quad\text{as }n\to\infty.$
@@
$U_{n}=\mathbb{E}[X_{R}(A)|\{Z_{i}\}]=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}\mathbb{E}[{\mathbb{I}}(A_{[i_{1},\ldots,i_{r}]}\cong R)|\{Z_{i}\}]=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}h(Z_{i_{1}},\ldots,Z_{i_{r}})$
@@
$\tilde{U}_{n}=\mathbb{E}[X_{R}(A)|\{\tilde{Z}_{i}\}]=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}h(\tilde{Z}_{i_{1}},\ldots,\tilde{Z}_{i_{r}}).$
@@
$).\displaystyle X_{R}(A)-X_{R}(\tilde{A})=(X_{R}(A)-U_{n})+(U_{n}-\tilde{U}_{n})+(\tilde{U}_{n}-X_{R}(\tilde{A}$)).
Conditioning on $\{Z_{i}\}$ and $\{\tilde{Z}_{i}\}$, respectively, the first and third term vanish in probability from Lemma E.4. By Theorem 3, we know that the empirical CDF of $\tilde{Z}_{i}$’s satisfies the requirement of Lemma E.2. And then we call Lemma E.3, and the second term also vanishes in probability. ∎